# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v2

@@ Abstract
Large Language Models (LLMs) struggle with hallucinations and outdated knowledge due to their reliance on static training data. Retrieval-Augmented Generation (RAG) mitigates these issues by integrating external dynamic information enhancing factual and updated grounding. Recent advances in multimodal learning have led to the development of Multimodal RAG, incorporating multiple modalities such as text, images, audio, and video to enhance the generated outputs. However, cross-modal alignment and reasoning introduce unique challenges to Multimodal RAG, distinguishing it from traditional unimodal RAG. This survey offers a structured and comprehensive analysis of Multimodal RAG systems, covering datasets, metrics, benchmarks, evaluation, methodologies, and innovations in retrieval, fusion, augmentation, and generation. We precisely review training strategies, robustness enhancements, and loss functions, while also exploring the diverse Multimodal RAG scenarios. Furthermore, we discuss open challenges and future research directions to support advancements in this evolving field. This survey lays the foundation for developing more capable and reliable AI systems that effectively leverage multimodal dynamic external knowledge bases. Resources are available at https://github.com/llm-lab-org/Multimodal-RAG-Survey.
Ask in Any Modality:
@@ 1 Introduction & Background
In recent years, significant breakthroughs have been achieved in language models, driven primarily by the advent of transformers Vaswani et al. (2017), enhanced computational capabilities, and the availability of large-scale training data Naveed et al. (2024). The emergence of foundational Large Language Models (LLMs) Ouyang et al. (2022); Grattafiori et al. (2024); Touvron et al. (2023); Qwen et al. (2025); Anil et al. (2023) has revolutionized natural language processing (NLP), demonstrating unprecedented capabilities in a wide range of tasks including instruction following Qin et al. (2024), sophisticated reasoning Wei et al. (2024c), In-context Learning Brown et al. (2020), and multilingual machine translation Zhu et al. (2024a). These advancements have elevated the performance of various NLP tasks, opening new avenues for research and application. Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al. (2024); Xu et al. (2024b). Their reliance on parametric memory restricts their ability to access up-to-date knowledge, making them less effective for knowledge-intensive tasks compared to task-specific architectures. Moreover, providing provenance for their decisions and updating their world knowledge remain critical open problems Lewis et al. (2020).
@@
RAG Lewis et al. (2020) has emerged as a promising solution to these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al. (2021); Ding et al. (2024a). By dynamically accessing vast external knowledge repositories, RAG systems enhance knowledge-intensive tasks while ensuring responses remain grounded in verifiable sources Gao et al. (2023).
In practice, RAG systems operate through a retriever-generator pipeline. The retriever leverages embedding models Chen et al. (2024a); Rau et al. (2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al. (2024a). These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses. Recent advancements in RAG frameworks, such as planning-guided retrieval Lee et al. (2024), agentic RAG An et al. (2024), and feedback-driven iterative refinement Liu et al. (2024c); Asai et al. (2023), further enhance both retrieval and generation stages. However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
@@
Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information. The introduction of CLIP (Contrastive Language-Image Pretraining) Radford et al. (2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al. (2024); Wang et al. (2023); Pramanick et al. (2023).
These breakthroughs have driven progress in various domains, including sentiment analysis Das and Singh (2023) and cutting-edge biomedical research Hemker et al. (2024), demonstrating the value of multimodal approaches. By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning plays a key role in advancing artificial general intelligence (AGI) Song et al. (2025).
Figure 1: Overview of the multimodal retrieval-augmented generation (RAG) pipeline, highlighting the advancements and techniques employed at each stage. The flow begins with query preprocessing, where user queries are refined and then encoded into a shared embedding space alongside a multimodal database. Retrieval strategies, such as modality-centric retrieval, similarity search, and re-ranking, enhance document selection, while fusion mechanisms align and integrate data from multiple modalities using score fusion or attention-based methods. Augmentation techniques such as iterative retrieval with feedback mechanisms, further refine the retrieved documents for the multimodal LLM. The generation stage incorporates innovations like Chain-of-Thought reasoning and source attribution for better outputs, with loss functions combining alignment loss and generation loss to optimize both retrieval and generation components. Noise management techniques are also applied to improve training stability and robustness.
@@
The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al. (2023a); Team et al. (2024); Li et al. (2023b). For example, GPT-4 OpenAI et al. (2024) demonstrates human-level performance in various benchmarks by processing both text and images, marking a significant milestone in multimodal perception and interaction. Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al. (2023); Chen et al. (2022a). This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs. However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al. (2023a). Figure 1 illustrates the general pipeline of these systems.
Task Formulation
A mathematical formulation of the general task for multimodal RAG is presented in this section. These systems generate a multimodal response, denoted as $r$, in response to a multimodal query $q$.
Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a multimodal corpus. Each document $d_{i}\in D$ is associated with a modality $M_{d_{i}}$ and processed by a modality-specific encoder, yielding:
$z_{i}=Enc_{M_{d_{i}}}(d_{i})$ (1)
The set of all encoded representations is denoted by $Z=\{z_{1},z_{2},...,z_{n}\}$. Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment. A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$. To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
$X=\{d_{i}\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\}$ (2)
where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations. The generative model $G$ produces the final multimodal response, given the user query $q$ and the retrieved documents $X$ as context, denoted as $r=G(q,X)$.
@@
As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems. While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al. (2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs. The only related survey to date Zhao et al. (2023a) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities. However, our survey provides a more detailed and innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth. Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community. In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
@@
In this work, (i) we provide a comprehensive review of multimodal RAG, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations in retrieval, fusion, augmentation, generation, training strategies, and loss functions. (ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers. (iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available. (iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
2 Datasets and Benchmarks
Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources. Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al. (2014), Flickr30K Young et al. (2014), and LAION-400M Schuhmann et al. (2021), while visual question answering with external knowledge is supported by OK-VQA Marino et al. (2019) and WebQA Chang et al. (2022). For complex multimodal reasoning, MultimodalQA Talmor et al. (2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al. (2015) and YouCook2 Zhou et al. (2018).
In the medical domain, MIMIC-CXR Johnson et al. (2019) and CheXpert Irvin et al. (2019) facilitate tasks such as medical report generation. It is noteworthy that a number of these datasets are unimodal (e.g., solely text-based or image-based). Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities. This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval. The $M^{2}RAG$ Ma et al. (2024c) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements Vision-focused evaluations, including MRAG-Bench Hu et al. (2024c), VQAv2 Goyal et al. (2017a) and VisDoMBench Suri et al. (2024), test models on complex visual tasks. Dyn-VQA Li et al. (2024b), MMBench Liu et al. (2025), and ScienceQA Saikh et al. (2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs.
Knowledge-intensive benchmarks, such as TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al. (2024), measure integration of unstructured and structured data. Advanced retrieval benchmarks such as RAG-Check Mortaheb et al. (2025a) evaluate retrieval relevance and system reliability, while specialized assessments such as Counterfactual VQA Niu et al. (2021) test robustness against adversarial inputs. Additionally, OCR impact studies such as OHRBench Zhang et al. (2024d) examine the cascading effects of errors on RAG systems. Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix (§B).
@@ 2 Datasets and Benchmarks
 (§4.1.1), fill=red!15 [ Maximum Inner Product Search (MIPS), fill=blue!15 [ TPU-KNN Chern et al. (2022), ScaNN Guo et al. (2020), MAXSIM score Chan and Ng (2008), ADQ Li et al. (2024a), Zhang et al. (2023a), BanditMIPS Tiwari et al. (2024), MUST Wang et al. (2024c), FARGO Zhao et al. (2023b), MuRAG Chen et al. (2022a), RA-CM3 Yasunaga et al. (2023), Nguyen et al. (2024), Graph-based ANNs Zhang et al. (2024a), Zhai et al. (2023), Deeperimpact Basnet et al. (2024), RetrievalAttention Liu et al. (2024a), FactMM-RAG Sun et al. (2024b) , leaf, text width=53em ] ] [ Multimodal
@@
Encoders, fill=blue!15 [ CLIP Radford et al. (2021), BLIP Li et al. (2022a), MARVEL Zhou et al. (2024c), ALIGN Jia et al. (2021), FLAVA Singh et al. (2022), UniVL-DR Liu et al. (2023b), UniIR Wei et al. (2024a), GME Zhang et al. (2024i), VISTA Zhou et al. (2024b), ColPali Faysse et al. (2024), InternVideo Wang et al. (2022), Ovis Lu et al. (2024) Mi-RAG Adjali et al. (2024) , leaf, text width=53em ] ] ] [ Modality-Centric
@@
Retrieval (§4.1.2), fill=red!15 [ Text-Centric, fill=blue!15 [ Contriever Izacard et al. (2022), GTE Li et al. (2023c), Re-Imagen Chen et al. (2022b), BM25 Robertson and Zaragoza (2009), MiniLM Wang et al. (2020a), BGE-M3 Chen et al. (2024b), CapRet Shohan et al. (2024), OMG-QA Nan et al. (2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al. (2024b), RAFT Zhang et al. (2024h), CRAG Yan et al. (2024), M2-RAG Ma et al. (2024c) , leaf, text width=53em ] ] [ Vision-Centric, fill=blue!15 [ VQA4CIR Feng et al. (2023), Unifashion Zhao et al. (2024), Jang et al. Jang et al. (2024), Pic2word Saito et al. (2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al. (2023), Joshi et al. Joshi et al. (2024), VISA Ma et al. (2024b), ImgRet Shohan et al. (2024), EchoSight Yan and Xie (2024), Xue et al. Xue et al. (2024) , leaf, text width=53em ] ] [ Video-Centric, fill=blue!15 [ iRAG Arefeen et al. (2024), VideoRAG Ren et al. (2025), VideoRAG Jeong et al. (2025), T-Mass Wang et al. (2024b), MV-Adapter Jin et al. (2024), OmAgent Zhang et al. (2024e), CM2 Kim et al. (2024), Video-RAG Luo et al. (2024b), CTCH Shen et al. (2024), RTime Du et al. (2024), VideoMAE Tong et al. (2022), DrVideo Ma et al. (2024d) , leaf, text width=53em ] ] [ Document Retrieval, fill=blue!15 [ ColPali Faysse et al. (2024), ColQwen2 Wang et al. (2024d), M3DocVQA Cho et al. (2024), ViTLP Mao et al. (2024), DocLLM Wang et al. (2024a), CREAM Zhang et al. (2024b), mPLUG-DocOwl 1.5 Hu et al. (2024a), mPLUG-DocOwl 2 Hu et al. (2024b), VisDom Suri et al. (2024), DSE Ma et al. (2024a) , leaf, text width=53em ] ] ] [ Re-ranking
@@
Strategies (§4.1.3), fill=red!15 [ Optimized Example Selection, fill=blue!15 [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024a), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=53em ] ] [ Relevance Score
@@
Evaluation, fill=blue!15 [ RAG-Check Mortaheb et al. (2025a, b), UniRaG Zhi Lim et al. (2024), MR2AG Zhang et al. (2024g), LDRE Yang et al. (2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al. (2024), OMG-QA Nan et al. (2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al. (2024a) , leaf, text width=53em ] ] [ Filtering
@@
and Alignment
 (§4.2.1), fill=red!15 [ M3 Cai et al. (2025)  Zhi Lim et al. (2024),  Sharifymoghaddam et al. (2024), REVEAL Hu et al. (2023), RAG-Driver Yuan et al. (2024), C3Net Zhang et al. (2024c), LLM-RA Jian et al. (2024),  Riedler and Langer (2024), VISA Ma et al. (2024b), MA-LMM He et al. (2024),  Xue et al. (2024), RA-BLIP Ding et al. (2024b), Re-IMAGEN Chen et al. (2022b), MegaPairs Zhou et al. (2024a), Wiki-LLaVA Caffagni et al. (2024), VISRAG Yu et al. (2024) , leaf, text width=62.0em ] ] [ Attention-Based
Mechanisms (§4.2.2), fill=red!15 [ RAMM Yuan et al. (2023), EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), RAGTrans Cheng et al. (2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al. (2024),  Xu et al. (2024a),  Kim et al. (2024), M2-RAAP Dong et al. (2024b), Mu-RAG Chen et al. (2022a), Ou et al. Ou et al. (2025), CADMR Khalafaoui et al. (2024) , leaf, text width=62.0em ] ] [ Unified Frameworks
and Projections (§4.2.3), fill=red!15 [ Hybrid-RAG Su et al. (2024a), Dense2Sparse Nguyen et al. (2024), IRAMIG Liu et al. (2024b), M3DocRAG Cho et al. (2024), DQU-CIR Wen et al. (2024), PDF-MVQA Ding et al. (2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al. (2024),  Li et al. (2022a) , leaf, text width=62.0em ] ] ] [ Augmentation

### v3

# A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

@@ Abstract
Large Language Models (LLMs) suffer from hallucinations and outdated knowledge due to their reliance on static training data. Retrieval-Augmented Generation (RAG) mitigates these issues by integrating external dynamic information for improved factual grounding. With advances in multimodal learning, Multimodal RAG extends this approach by incorporating multiple modalities such as text, images, audio, and video to enhance the generated outputs. However, cross-modal alignment and reasoning introduce unique challenges beyond those in unimodal RAG. This survey offers a structured and comprehensive analysis of Multimodal RAG systems, covering datasets, benchmarks, metrics, evaluation, methodologies, and innovations in retrieval, fusion, augmentation, and generation. We review training strategies, robustness enhancements, loss functions, and agent-based approaches, while also exploring the diverse Multimodal RAG scenarios. In addition, we outline open challenges and future directions to guide research in this evolving field. This survey lays the foundation for developing more capable and reliable AI systems that effectively leverage multimodal dynamic external knowledge bases. All resources are publicly available  [^1] https://github.com/llm-lab-org/Multimodal-RAG-Survey.
spacing=nonfrench
Ask in Any Modality
@@ 1 Introduction & Background
Recent advancements in transformer architectures Vaswani et al. (2017), coupled with increased computational resources and the availability of large-scale training datasets Naveed et al. (2024), have significantly accelerated progress in the development of language models. The emergence of foundational Large Language Models (LLMs) Ouyang et al. (2022); Grattafiori et al. (2024); Touvron et al. (2023); Qwen et al. (2025); Anil et al. (2023), has revolutionized natural language processing (NLP), excelling in tasks such as instruction following Qin et al. (2024), reasoning Wei et al. (2024b), in-context learning Brown et al. (2020), and multilingual translation Zhu et al. (2024a). Despite these achievements, LLMs face challenges such as hallucinations, outdated knowledge, and a lack of verifiable reasoning Huang et al. (2024); Xu et al. (2024b). Their reliance on parametric memory limits access to up-to-date information, reducing their effectiveness in knowledge-intensive tasks.
@@
RAG Lewis et al. (2020) addresses these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al. (2021); Ding et al. (2024a). By dynamically accessing external knowledge sources, RAG enhances knowledge-intensive tasks while grounding responses in verifiable sources Gao et al. (2023). In practice, RAG systems follow a retriever-generator pipeline: the retriever uses embedding models Chen et al. (2024a); Rau et al. (2024) to identify relevant passages from external knowledge bases and may apply re-ranking techniques to improve precision Dong et al. (2024a). The retrieved passages are then provided to the generator, which leverages this contextual information to produce more informed and coherent responses. Recent advancements in RAG frameworks, such as planning-guided retrieval Lee et al. (2024), agentic RAG An et al. (2024), and feedback-driven iterative refinement Liu et al. (2024c); Asai et al. (2023), have further improved both the retrieval and generation components of these systems.
Figure 1: Overview of the multimodal RAG pipeline, illustrating key techniques and recent advancements.
@@
In parallel with advances in language modeling, multimodal learning has emerged as a transformative area in artificial intelligence, enabling systems to integrate and reason over heterogeneous data sources for more comprehensive representations. A pivotal breakthrough was the introduction of CLIP model Radford et al. (2021), which aligned visual and textual modalities through contrastive learning and inspired a wave of subsequent models Alayrac et al. (2024); Wang et al. (2023a); Pramanick et al. (2023). These developments have catalyzed progress across diverse domains, including sentiment analysis Das and Singh (2023) and biomedical applications Hemker et al. (2024), highlighting the effectiveness of multimodal approaches. By facilitating the joint processing of text, images, audio, and video, multimodal learning is increasingly recognized as a critical enabler of progress toward artificial general intelligence (AGI) Song et al. (2025).
@@
The extension of large language models to multimodal LLMs (MLLMs) has significantly broadened their capabilities, enabling reasoning and generation across multiple data modalities Liu et al. (2023a); Team et al. (2024); Li et al. (2023b). Notably, GPT-4 OpenAI et al. (2024) demonstrates human-level performance by jointly processing text and images, marking a milestone in multimodal understanding. Building on this progress, multimodal RAG incorporates diverse sources, such as images, audio, and structured data, to enrich contextual grounding and enhance generation quality Hu et al. (2023); Chen et al. (2022a). This approach improves reasoning by leveraging cross-modal cues, but also introduces challenges, including modality selection, effective fusion, and managing cross-modal relevance Zhao et al. (2023a). Figure 1 illustrates the general pipeline of these systems.
Multimodal RAG Formulation
We present a mathematical formulation of multimodal RAG. These systems aim to generate a multimodal response $r$ given a multimodal query $q$. Let $D=\{d_{1},d_{2},\ldots,d_{n}\}$ denote a multimodal corpus. For clarity, we assume each document $d_{i}\in D$ is associated with a single modality $M_{d_{i}}$. In practice, however, documents often span multiple modalities—for example, a scientific article containing both text and images. Such cases are typically addressed by either decomposing the document into modality-specific sub-documents or employing universal encoders capable of jointly processing multiple modalities.
Each document $d_{i}$ is encoded using its corresponding modality-specific encoder, yielding $z_{i}=\text{Enc}_{M_{d_{i}}}(d_{i})$. The collection of all encoded representations is denoted as $Z=\{z_{1},z_{2},\ldots,z_{n}\}$. These modality-specific encoders project diverse input modalities into a shared semantic space, enabling cross-modal alignment.
A retrieval model $R$ computes a relevance score $s(e_{q},z_{i})$ between the encoded query representation $e_{q}$ (obtained by encoding $q$ using the appropriate encoders) and each document representation $z_{i}$. The retrieval-augmented multimodal context $X$ is constructed by selecting documents whose relevance scores exceed a modality-specific threshold:
$X=\{d_{i}\in D\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\},\vspace{-0.1cm}$
where $\tau_{M_{d_{i}}}$ is the relevance threshold for the modality $M_{d_{i}}$, and $s$ is the scoring function that measures semantic relevance. Finally, the generative model $G$ produces the response conditioned on the original query $q$ and the retrieved context $X$, formally defined as $r=G(q,X)$.
@@
Multimodal RAG is a rapidly emerging field, yet a comprehensive survey dedicated to its recent advancements remains lacking. While over ten surveys discuss RAG-related topics such as Agentic RAG Singh et al. (2025), none specifically focus on the multimodal setting. To our knowledge, the only relevant work Zhao et al. (2023a) categorizes multimodal RAGs by application and modality. In contrast, our survey adopts a more innovation-driven perspective, offering a detailed taxonomy and addressing recent trends and open challenges. We review over 100 recent papers, primarily from the ACL Anthology, reflecting the growing interest and progress in this domain.
@@
In this work, (i) we present a comprehensive review of multimodal RAG, covering task formulation, datasets, benchmarks, applications, and key innovations across retrieval, fusion, augmentation, generation, training strategies, loss functions, and agent frameworks. (ii) We propose a structured taxonomy (Figure 2) that categorizes state-of-the-art models by their core contributions, highlighting methodological advances and emerging trends. (iii) We provide open-access resources, including datasets, benchmarks, and implementation details, to facilitate future research. (iv) Finally, we identify research gaps and offer insights to guide future directions in this rapidly evolving field.
2 Datasets, Evaluation, and Applications
We review diverse datasets and benchmarks supporting tasks such as multimodal summarization, visual QA, video understanding, and more. For full details, see Appendix (§B) and Tables 1 and 2. Multimodal RAG has been applied across various domains, including healthcare, software engineering, fashion, entertainment, and emerging fields. An overview of tasks and applications are detailed in Appendix (§E) and  Figure 3. Evaluating these systems requires multiple metrics, covering retrieval performance, generation quality, and modality alignment. The complete evaluation methods, metrics, and their definitions and formulations are in Appendix (§C).
@@ 2 Datasets, Evaluation, and Applications
 (§3.1), fill=red!15 [ Maximum Inner Product Search (MIPS), fill=blue!15 [ TPU-KNN Chern et al. (2022), ScaNN Guo et al. (2020), MAXSIM score Chan and Ng (2008), ADQ Li et al. (2024a), Zhang et al. (2023a), BanditMIPS Tiwari et al. (2024), MUST Wang et al. (2024c), FARGO Zhao et al. (2023b), MuRAG Chen et al. (2022a), RA-CM3 Yasunaga et al. (2023), Nguyen et al. (2024), Graph-based ANNs Zhang et al. (2024a), Zhai et al. (2023), Deeperimpact Basnet et al. (2024), RetrievalAttention Liu et al. (2024a), FactMM-RAG Sun et al. (2024b) , leaf, text width=53em ] ] [ Multimodal
@@
Encoders, fill=blue!15 [ CLIP Radford et al. (2021), BLIP Li et al. (2022), MARVEL Zhou et al. (2024c), ALIGN Jia et al. (2021), FLAVA Singh et al. (2022), UniVL-DR Liu et al. (2023b), UniIR Wei et al. (2024a), GME Zhang et al. (2024i), VISTA Zhou et al. (2024b), ColPali Faysse et al. (2024), InternVideo Wang et al. (2022), Ovis Lu et al. (2024), LLaVE Lan et al. (2025) Mi-RAG Adjali et al. (2024) , leaf, text width=53em ] ] ] [ Modality-Centric
@@
Retrieval (§3.1), fill=red!15 [ Text-Centric, fill=blue!15 [ Contriever Izacard et al. (2022), GTE Li et al. (2023c), Re-Imagen Chen et al. (2022b), BM25 Robertson and Zaragoza (2009), MiniLM Wang et al. (2020a), BGE-M3 Chen et al. (2024b), CapRet Shohan et al. (2024), OMG-QA Nan et al. (2024b), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al. (2024b), RAFT Zhang et al. (2024h), CRAG Yan et al. (2024), M2-RAG Ma et al. (2024d) , leaf, text width=53em ] ] [ Vision-Centric, fill=blue!15 [ VQA4CIR Feng et al. (2023), Unifashion Zhao et al. (2024), Jang et al. Jang et al. (2024), Pic2word Saito et al. (2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al. (2023), Joshi et al. Joshi et al. (2024), VISA Ma et al. (2024b), ImgRet Shohan et al. (2024), EchoSight Yan and Xie (2024), Xue et al. Xue et al. (2024b) , leaf, text width=53em ] ] [ Video-Centric, fill=blue!15 [ iRAG Arefeen et al. (2024), VideoRAG Ren et al. (2025), VideoRAG Jeong et al. (2025), T-Mass Wang et al. (2024b), MV-Adapter Jin et al. (2024), OmAgent Zhang et al. (2024e), CM2 Kim et al. (2024), Video-RAG Luo et al. (2024b), CTCH Shen et al. (2024), RTime Du et al. (2024), VideoMAE Tong et al. (2022), DrVideo Ma et al. (2024e) , leaf, text width=53em ] ] [ Audio-Centric, fill=blue!15 [ CA-CLAP Xue et al. (2024a), Recap Ghosh et al. (2024), SpeechRAG Min et al. (2025), WavRAG Chen et al. (2025b), SEAL Sun et al. (2025), Audiobox TTA-RAG Yang et al. (2024a), DRCap Li et al. (2025c), P2PCAP Changin et al. (2024), LA-RAG Li et al. (2024b), Xiao et al. (2025) , leaf, text width=53em ] ] [ Document
Retrieval, fill=blue!15 [ ColPali Faysse et al. (2024), ColQwen2 Wang et al. (2024d), M3DocVQA Cho et al. (2024), ViTLP Mao et al. (2024), DocLLM Wang et al. (2024a), CREAM Zhang et al. (2024b), mPLUG-DocOwl 1.5 Hu et al. (2024a), mPLUG-DocOwl 2 Hu et al. (2024b), VisDom Suri et al. (2024), DSE Ma et al. (2024a) , SV-RAG Chen et al. (2025a) , leaf, text width=53em ] ] ] [ Re-ranking
Strategies (§3.1), fill=red!15 [ Optimized
Example
@@
Selection, fill=blue!15 [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024a), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=53em ] ] [ Relevance Score
@@
Evaluation, fill=blue!15 [ RAG-Check Mortaheb et al. (2025a, b), UniRaG Zhi Lim et al. (2024), MR2AG Zhang et al. (2024g), LDRE Yang et al. (2024b), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al. (2024), OMG-QA Nan et al. (2024b), EchoSight Yan and Xie (2024), EgoInstructor Xu et al. (2024a), VR-RAG Khan et al. (2025) , leaf, text width=53em ] ] [ Filtering
@@
and Alignment (§3.2), fill=red!15 [ M3 Cai et al. (2025)  Zhi Lim et al. (2024),  Sharifymoghaddam et al. (2024), REVEAL Hu et al. (2023), RAG-Driver Yuan et al. (2024), C3Net Zhang et al. (2024c), LLM-RA Jian et al. (2024),  Riedler and Langer (2024), VISA Ma et al. (2024b), MA-LMM He et al. (2024),  Xue et al. (2024b), RA-BLIP Ding et al. (2024b), Re-IMAGEN Chen et al. (2022b), MegaPairs Zhou et al. (2024a), Wiki-LLaVA Caffagni et al. (2024), VISRAG Yu et al. (2024) , leaf, text width=62.0em ] ] [ Attention-Based
Mechanisms (§3.2), fill=red!15 [ RAMM Yuan et al. (2023), EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), RAGTrans Cheng et al. (2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al. (2024),  Xu et al. (2024a),  Kim et al. (2024), M2-RAAP Dong et al. (2024b), Mu-RAG Chen et al. (2022a), Ou et al. Ou et al. (2025), CADMR Khalafaoui et al. (2024) , leaf, text width=62.0em ] ] [ Unified Frameworks
and Projections (§3.2), fill=red!15 [ Hybrid-RAG Su et al. (2024a), Dense2Sparse Nguyen et al. (2024), IRAMIG Liu et al. (2024b), M3DocRAG Cho et al. (2024), DQU-CIR Wen et al. (2024), PDF-MVQA Ding et al. (2024d), SAM-RAG Zhai (2024), UFineBench Zuo et al. (2024),  Li et al. (2022) , leaf, text width=62.0em ] ] ] [ Augmentation